a car has an initial velocity of 30m/s and an acceleration of -5 m/s2. Find its displacement after 10 seconds.



1
Expert's answer
2021-09-27T09:02:05-0400

We use the information provided and we find the displacement for a movement with linear acceleration:


"d_0=0\,m
\\ a=-5\frac{m}{s^2}
\\ v_i=30\frac{m}{s}
\\t=10\,s
\\d=d_0+v_0t+\cfrac{at^2}{2}
\\d=0\,m+(30\frac{m}{s})(10\,s)+\cfrac{-5\frac{m}{s^2}(10\,s)^2}{2}
\\ \implies d=0\,m+300\,m-250\,m=50\,m"


After substitution, we find that the displacement of the car after 10 seconds is about 50 m.



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.